OpenStudy (howard-wolowitz):

P(X<2)..... using the table

OpenStudy (howard-wolowitz):

@Directrix

Directrix (directrix):

The probability of X being less than or equal to 2 is: P(1) + P(2) = 1/6 + 1/6 = ? @Howard-Wolowitz

OpenStudy (howard-wolowitz):

i got 0.333333333333

Directrix (directrix):

I got 1/6 + 1/6 = 2/6 = 1/3.